Please indicate the correct answer and explain. Thank you!

Kali is a dot-com entrepreneur who has established a Web site at which people can design and buy aring. Kali pays $600 a month for a Web server and Internet connection. The rings that customers design are made to order by another firm, and Kali pays this firm $40 a ring. Kali has no other costs The table shows the demand schedule for Kalis rings Price dollars per ring) 200 160 120 80 40 Quantity (rings per month) 40 80 120 160 200 Kalis profit-maximizing output is □ rings a month. Kalis profit-maximizing price is aring. Kalis economic profit is S□ a month.
Answer #1

饒 TR MR MC 2 C) 40 6400 80 9600 120 9600 160 6400 0 160 120 80 40 2 40 200 0160

Since fixed cost is $600 per month for internet connection and $40 per ring is a marginal cost.

Since the profit-maximizing condition is

MR=MC

Hence corresponding to this condition the profit-maximzing price is $120.

Kali's price=$120

Kali's Quantity= 80 units

Kali's economic profit= TR - TC

=9600- (FC+MC*Q)

=9600 -(600+40*80)

=9600- (600+3200)

=9600-3800

=$5,800